### v8

@@ 3.4 Cross-Stage Dual-Domain Multi-Scale Attention Module
In transmission line defect detection, most objects are small-scale objects such as cracks, rust spots, or insulator damages. These defects often exhibit weak visual saliency and subtle structural variations, making them difficult to distinguish from complex backgrounds like vegetation, sky, or towers. Conventional convolutional networks are inherently constrained by their limited local receptive fields, and simply enlarging the kernel size to capture global context quickly leads to over-parameterization and substantial computational overhead, while also causing oversmoothing and the loss of fine-grained detailsFinder et al. (2025). Similarly, traditional attention mechanisms that focus solely on either spatial or channel domains fail to provide sufficient enhancement for ambiguous and small-scale objectsZhou et al. (2025). Therefore, a more effective mechanism is required to simultaneously preserve detailed structural information, capture global semantic context, and adaptively emphasize defect-related features.
After an extensive literature review, we find that coupling frequency decomposition with attention is particularly effective for tiny, edge-dominated defects in UAV imagery—preserving high-frequency details and suppressing low-frequency clutter that misleads purely spatial attention. Therefore, we propose CSDMAM, which injects learnable frequency gating directly into the attention formation while retaining anisotropic, large-kernel spatial paths tailored to line-shaped defects.
To improve the effectiveness of transmission lines defect detection, we propose the Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM). Given an input feature map $X\in\mathbb{R}^{B\times C\times H\times W}$, a $1\times 1$ convolution is first applied to adjust the channel dimension. The transformed features are then processed in parallel through multiple specialized branches, and their outputs are aggregated via residual summation before being projected by another $1\times 1$ convolution. This design enables the module to capture diverse feature representations while maintaining computational efficiency. The overall architecture of CSDMAM is illustrated in Figure 5.
@@
Specifically, to enlarge the receptive field, we add a large-receptive-field branch employing depth-wise convolutions with kernel sizes $(k,k),(1,k),(k,1)$, where $k=31$. The square kernel models global contextual information, while the strip kernels capture elongated dependencies along horizontal and vertical directions, which are especially important for line-shaped structures in transmission infrastructure.
To preserve fine-grained local cues, we further introduce a local branch with a lightweight depth-wise $1\times 1$ convolution, ensuring that subtle defects such as cracks or corrosion spots are not smoothed out by large-kernel operations.
To strengthen semantic discrimination and capture long-range contextual information, we design a Global Branch built upon the Frequency–Spatial–Channel Attention (FSCA) module, which jointly exploits frequency, spatial, and channel dependencies. Given the input features $X\in\mathbb{R}^{B\times C\times H\times W}$, a two-dimensional Fourier transform is first applied to obtain their frequency-domain representation:
@@
where $\mathcal{F}$ denotes the discrete Fourier transform (DFT).
Channel-wise weights are then computed via global average pooling followed by a $1\times 1$ convolution, producing $A_{c}\in\mathbb{R}^{B\times C\times 1}$. These weights are used to enhance the frequency representation:
@@
which is subsequently transformed back into the spatial domain using the inverse DFT:
@@
Next, a spatial–channel reweighting map $A_{sc}$ is generated through global pooling and a $1\times 1$ convolution. This map adaptively emphasizes defect-relevant regions and further refines the feature representation, resulting in the final output of the Global Branch:
@@
To effectively integrate the complementary information captured by all branches, we introduce a residual aggregation branch. In this step, the outputs of the large-receptive-field branch, the local branch, and the FSCA branch are combined with the original input through element-wise summation. This residual fusion mechanism not only preserves the original feature information but also enhances gradient flow and stabilizes training. By consolidating global context, fine-grained details, and dual-domain attention responses into a unified representation, the residual aggregation branch ensures comprehensive feature enhancement and precise defect localization. The final aggregated response is formulated as
@@
where $DW$ denotes depth-wise convolution with the specified kernel size.
Finally, to further improve computational efficiency and facilitate gradient propagation, we extend the design into a cross-stage structure. Following the Cross Stage Partial (CSP) strategy, the input is split into two parts: one processed by the multi-branch attention module and the other preserved as an identity mapping. The two parts are then concatenated and fused using a $1\times 1$ convolution to produce the final output:
$Y_{cs}=\text{Conv}\left([\text{DMAM}(X_{\text{branch}}),X_{\text{identity}}]\right).$
By jointly leveraging large receptive fields, local detail preservation, FSCA-based dual-domain enhancement, and cross-stage fusion, CSDMAM achieves a favorable balance between global semantic understanding and precise defect localization. This design significantly improves the detection of small, visually ambiguous defects in UAV inspection imagery.
@@ 3.5.2 Focaler-IoU Overlap Term
3.5.3 Angle-Weighted Distance Penalty (SIoU-style)
Let $\phi\in[0,\pi/4]$ be the acute angle between $\Delta$ and its nearest coordinate axis:
@@ 3.5.3 Angle-Weighted Distance Penalty (SIoU-style)
where $10^{-4}$ prevents division by zero when $\Delta\to 0$. The angle cost in SIoU can be written as $\Lambda=1-2\sin^{2}\!\bigl(\phi-\tfrac{\pi}{4}\bigr)$. Using $\cos(2\theta)=1-2\sin^{2}\theta$ and $\cos(2\phi-\tfrac{\pi}{2})=\sin(2\phi)$, we obtain the exact equivalence
$\Lambda=\sin(2\phi).$ (8)
We then define
@@
so that the distance term below naturally takes the canonical form $1-\exp\bigl((2-\Lambda)\cdot(\cdot)\bigr)$: larger misalignment (large $2-\Lambda$) increases the penalty before purely shrinking the radial offset.
We penalize center offsets per axis with squared, scale-normalized terms:
$\mathrm{Dist}=2-\exp\!\Bigl(\mathrm{Angle}\cdot\bigl(\tfrac{d_{x}}{w_{\mathrm{box}}}\bigr)^{\!2}\Bigr)-\exp\!\Bigl(\mathrm{Angle}\cdot\bigl(\tfrac{d_{y}}{h_{\mathrm{box}}}\bigr)^{\!2}\Bigr).$ (10)
Because $\mathrm{Angle}\leq 0$ by Eq. (9), each exponential factor decreases as the (normalized) offset grows, hence $\mathrm{Dist}$ is monotone in misalignment while remaining smooth around zero. Axis-wise normalization ensures scale invariance across object sizes, contributing to the overall scale invariance of the loss.
@@ 3.5.6 Wise Difficulty Modulation via EMA Normalization
We quantify sample difficulty with the normalized IoU gap
$\beta=\frac{1-\mathrm{IoU}}{\mathbb{E}[\,1-\mathrm{IoU}\,]},\qquad\mathbb{E}[\,1-\mathrm{IoU}\,]\leftarrow(1-m)\,\mathbb{E}[\,1-\mathrm{IoU}\,]+m\cdot(1-\mathrm{IoU})_{\text{batch}},$ (13)
where the expectation is tracked by an exponential moving average (EMA). We use $m=10^{-2}$, which is stable yet responsive for shifts across training phases; in practice we add a tiny $\varepsilon=10^{-6}$ to the denominator for numerical stability. Crucially, Eq. (13) operates on the raw IoU from Eq. (3) to preserve geometric interpretability.
@@
This non-monotonic weighting targets informative moderately hard samples around $\beta\!\approx\!\delta$, gently down-weights trivially easy cases ($\beta\!\ll\!\delta$) and prevents domination by extreme outliers as $\beta\!\gg\!\delta$.
@@ 3.5.7 Final Objective and Batch-wise Computation
Multiplying the geometric core (Eq. (12)) by the difficulty weight (Eq. (14)) yields the Focaler–Wise–SIoU objective:
$\boxed{L\;=\;\Bigl(1-\hat{\mathrm{IoU}}+\tfrac{1}{2}(\mathrm{Dist}+\mathrm{Shape})\Bigr)\cdot\gamma(\beta),\quad\beta=\frac{1-\mathrm{IoU}}{\mathbb{E}[\,1-\mathrm{IoU}\,]}}$ (15)
We now detail a concise batch-wise procedure that operationalizes Eqs. (5)–(11) and (13)–(14) into a practical training routine:
Algorithm 1 Per-mini-batch computation of Focaler–Wise–SIoU
1:predictions $B_{p}$, targets $B_{t}$; running mean $\mu=\mathbb{E}[1-\mathrm{IoU}]$; hyperparameters $(d,u)=(0,0.95)$, $m=10^{-2}$, $\theta=4$, $(\alpha,\delta)=(1.7,2.7)$; $\varepsilon=10^{-6}$
2:Raw/Proxy IoU: compute raw $\mathrm{IoU}$ (Eq. (3)); compute $\hat{\mathrm{IoU}}$ (Eq. (5)); set $L_{\mathrm{IoU}}=1-\hat{\mathrm{IoU}}$ (Eq. (6))
3:Distance geometry: compute $\phi$ (Eq. (7)), $\mathrm{Angle}$ (Eq. (9)), $w_{\mathrm{box}},h_{\mathrm{box}}$ (Eq. (4)); then $\mathrm{Dist}$ (Eq. (10))
4:Shape: compute $\mathrm{Shape}$ (Eq. (11)) with $\theta$
5:Core loss: $L_{\mathrm{SIoU}}\leftarrow L_{\mathrm{IoU}}+\tfrac{1}{2}(\mathrm{Dist}+\mathrm{Shape})$ (Eq. (12))
6:Difficulty (raw IoU): $\beta\leftarrow(1-\mathrm{IoU})/(\mu+\varepsilon)$ (Eq. (13)); $w\leftarrow\gamma(\beta)=\beta/(\delta\,\alpha^{(\beta-\delta)})$ (Eq. (14))
7:Aggregate: $L\leftarrow\mathrm{mean}\bigl(L_{\mathrm{SIoU}}\cdot w\bigr)$ over positive matches (Eq. (15))
8:EMA update: $\mu\leftarrow(1-m)\mu+m\cdot\mathrm{mean}(1-\mathrm{IoU})$ (Eq. (13))
4 Experiments and Results Analysis
@@ 4.2 Model Training and Evaluation Metrics
We implement the proposed model on the PyTorch framework and build it upon the Ultralytics RT-DETR backbone initialized with pretrained weights. Training is conducted on a workstation with an NVIDIA GeForce RTX 3090 (24 GB) running Ubuntu 20.04, Python 3.10.14, PyTorch 2.3.1, and CUDA 12.1. The model is trained for 100 epochs with a batch size of 4 and an input resolution of $640\times 640$, using AdamW with an initial learning rate $1.0\times 10^{-4}$, a final learning-rate factor $lrf=1.0$, momentum 0.9, and weight decay $1.0\times 10^{-4}$. We apply a 2000-iteration warm-up with momentum set to 0.8 and a bias learning rate of 0.1.

### v9

@@ 1 Introduction
From a system perspective, the present work focuses on the post-processing stage of the inspection workflow, where UAVs execute routine patrols, capture high-resolution imagery along the transmission corridor, and subsequently upload these data to a ground-based server for automated analysis. In this paradigm, defect detection is conducted offline as a post-processing task rather than on-board the UAV, thus avoiding the stringent real-time and energy constraints associated with aerial platforms. This server-side processing setting enables the adoption of more expressive Transformer-based architectures and advanced optimization strategies while still meeting operational maintenance timelines. Meanwhile, the proposed model remains sufficiently lightweight to support potential migration to edge-computing platforms, achieving a practical balance between detection accuracy, robustness, and deployment flexibility.
@@ 3.4 Cross-Stage Dual-Domain Multi-Scale Attention Module
Transmission line defects in UAV imagery—such as cracks, rust spots, and insulator damages—are typically small, weakly salient, and embedded in cluttered backgrounds of vegetation, sky, or towers. Their ambiguous boundaries and subtle textural cues challenge conventional CNNs, whose local receptive fields hinder simultaneous modeling of global context and fine detail. Simply enlarging kernels inflates parameters and computation, often causing oversmoothing that erases critical structuresFinder et al. (2025). Likewise, attention mechanisms restricted to a single domain—spatial or channel—offer limited gains for tiny, low-contrast targetsZhou et al. (2025). Building on the joint frequency–spatial perspective outlined in Section 2.2, we introduce the Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM), which unifies frequency-domain selectivity with spatial-domain localization in an end-to-end attention framework.
The central idea is to embed learnable frequency gating directly within attention formation while retaining anisotropic, large-kernel spatial paths tailored to line-shaped defects. Frequency responses preserve edges, textures, and other high-frequency cues and suppress low-frequency clutter, whereas spatial aggregation localizes potential defects under complex context. Rather than treating frequency information as an auxiliary branch, CSDMAM integrates it adaptively so that frequency cues steer attention weights toward defect-relevant structures and away from distractors.
Given an input feature map $X\in\mathbb{R}^{B\times C\times H\times W}$, the module first applies a $1\times 1$ convolution to adjust channel dimensionality. The transformed tensor then flows through several parallel, specialized branches that operate across complementary scales. One branch performs a learnable frequency decomposition whose gated responses accentuate sharp structural variations indicative of defects while attenuating smooth backgrounds; these gated maps participate directly in attention computation. In parallel, an anisotropic large-kernel spatial path—implemented with directional and dilated filters—approximates a wide receptive field without excessive parameters and mitigates oversmoothing, thereby enhancing sensitivity to elongated conductor- and insulator-like patterns. Additional scale-selective paths aggregate broader semantic context and preserve fine detail. The branch outputs are merged by residual summation and projected with a final $1\times 1$ convolution, as illustrated in Figure 5.
Through this dual-domain, multi-scale design, CSDMAM preserves high-frequency, edge-dominated details, captures long-range semantics with modest overhead, and adaptively highlights subtle, defect-related responses. The resulting module addresses the principal failure modes of purely convolutional and single-domain attention approaches, delivering robust detection of small, low-saliency defects in complex aerial scenes.
@@
Large-Receptive-Field Branch. To enlarge the receptive field, we employ depthwise-separable convolutions with kernels $(k,k)$, $(1,k)$ and $(k,1)$, where $k=31$. The square $k\times k$ kernel aggregates global context, while the horizontal and vertical strip kernels capture elongated dependencies that are critical for the line-shaped structures typical of transmission infrastructure. The choice $k=31$ was determined by systematic empirical evaluation rather than by heuristic selection (see Section 4.3). Kernels substantially smaller than $31$ fail to capture the long-range, directionally elongated dependencies of line-like defects, whereas substantially larger kernels introduce excessive smoothing and disproportionate computational overhead without commensurate accuracy gains. Therefore, a $31\times 31$ depthwise-separable kernel provides the most favorable trade-off between global context aggregation and detail preservation for our task. We further restrict $k$ to the family $k=2^{n}-1$ (here $31=2^{5}-1$) for two practical reasons. First, an odd kernel size preserves a single central pixel, which simplifies symmetric, centered convolutional responses and padding. Second, when features are aggregated across hierarchical stages that expand spatial support by factors of two—via pooling, striding, or stage-wise downsampling—the geometric series $\sum_{i=0}^{n-1}2^{i}=2^{n}-1$ naturally arises. Selecting $k=2^{n}-1$ aligns the kernel’s spatial support with an $n$-level binary-scale expansion, facilitating efficient cross-scale context aggregation and predictable padding behavior. The ablation in Section 4.3 confirms that $k=31$ is optimal in our application.
Local Detail Preservation Branch. To prevent important fine-grained cues from being washed out by large-kernel operations, we introduce a lightweight local path using a depth-wise $1\times 1$ convolution, which preserves high-frequency details such as micro-cracks and corrosion spots while adding negligible computational burden.
Global Branch To strengthen semantic discrimination and capture long-range contextual information, we design a global branch built upon the FSCA module that jointly exploits frequency-, spatial-, and channel-wise dependencies. Given the input $X$, we first compute the frequency-domain representation
@@
where $\mathcal{F}$ denotes the discrete Fourier transform (DFT). Channel-wise weights are then obtained via global average pooling followed by a $1\times 1$ convolution, yielding $A_{c}\in\mathbb{R}^{B\times C\times 1}$. These weights modulate the frequency response,
@@
which is transformed back to the spatial domain through the inverse DFT,
@@
Finally, a spatial–channel reweighting map $A_{sc}$ is generated via global pooling and a $1\times 1$ convolution to adaptively emphasize defect-relevant regions, producing the branch output
@@
The effectiveness of FSCA and the integration of frequency-domain cues is validated in Section 4.3.
Residual Aggregation Branch. To effectively integrate complementary information from all branches while preserving the original signals, we adopt a residual fusion mechanism that combines the outputs of the large-receptive-field branch, the local branch, and the FSCA branch with the input via element-wise summation:
@@
where $DW$ denotes depth-wise convolution with the specified kernel size. This design enhances gradient flow, stabilizes optimization, and consolidates global context, fine details, and dual-domain attention responses into a unified representation for precise defect localization.
To further improve computational efficiency and facilitate gradient propagation, we extend CSDMAM into a cross-stage structure following the Cross Stage Partial (CSP) strategy. The input is split into two parts: one processed by the multi-branch attention module and the other preserved as an identity mapping. The two parts are concatenated and fused by a $1\times 1$ convolution to produce the final output:
$Y_{cs}=\mathrm{Conv}\big([\mathrm{DMAM}(X_{\text{branch}}),\,X_{\text{identity}}]\big).$
By jointly leveraging large receptive fields, local detail preservation, FSCA-based dual-domain enhancement, and cross-stage fusion, CSDMAM achieves a favorable balance between global semantic understanding and precise defect localization, significantly improving the detection of small, visually ambiguous defects in UAV inspection imagery.
@@ 3.5.2 Focaler-IoU Overlap Term
3.5.3 Angle-Weighted Distance Penalty
In transmission line defect detection tasks, objects such as insulator cracks, missing bolts, or bird nests often appear at arbitrary orientations and with substantial variation in scale and spatial distribution. These defects are typically small, elongated, and densely distributed along complex backgrounds such as towers or cables, especially in aerial inspection imagery. In such conditions, even minor center displacements between predicted and ground-truth bounding boxes can result in significant localization errors, particularly when the offset directions are not well aligned. Conventional distance-based penalties treat all directions uniformly—whether horizontal, vertical, or diagonal—thereby neglecting the geometric structure inherent in these domain-specific misalignments.
This limitation motivates us to develop an angle-weighted distance penalty that dynamically adjusts the localization penalty according to the angular alignment of the offset vector. By doing so, the model is encouraged to favor axis-aligned center corrections over diagonal ones, improving directional sensitivity and enhancing localization accuracy under the challenging settings encountered in power line defect detection.
We define the offset between the predicted and ground-truth box centers as $\Delta=(d_{x},d_{y})$, and let $\phi\in[0,\pi/4]$ denote the acute angle between $\Delta$ and its nearest coordinate axis. This angle reflects the degree of directional misalignment: $\phi=0$ indicates perfect alignment with one axis, while $\phi=\pi/4$ corresponds to a fully diagonal offset. It is computed as:
@@ 3.5.3 Angle-Weighted Distance Penalty
where a small constant in the denominator ensures numerical stability near zero displacement.
In the original SIoU formulation, the angle cost is defined as:
$\Lambda=1-2\sin^{2}\left(\phi-\tfrac{\pi}{4}\right)=\sin(2\phi),$
which increases as the offset becomes more axis-aligned. When the displacement vector approaches either the horizontal or vertical axis, the angle $\phi$ becomes small, resulting in a higher $\Lambda$, indicating better alignment between the predicted and target boxes.
To incorporate this angular alignment into the penalty formulation, we define a scaling factor that modulates the influence of the distance terms:
@@
which ensures that $\mathrm{Angle}\leq 0$. This negative-valued quantity is then used in an exponential decay to penalize displacement, such that larger directional misalignment results in steeper penalties. Specifically, when the offset direction is diagonal (i.e., small $\Lambda$), the penalty grows more aggressively; when it is axis-aligned (i.e., large $\Lambda$), the penalty grows more slowly. This behavior makes the model more sensitive to the type of misalignment rather than only its magnitude.
Based on this principle, we formulate the distance penalty as:
$\mathrm{Dist}=2-\exp\left(\mathrm{Angle}\cdot\left(\frac{d_{x}}{w_{\mathrm{box}}}\right)^{2}\right)-\exp\left(\mathrm{Angle}\cdot\left(\frac{d_{y}}{h_{\mathrm{box}}}\right)^{2}\right),$ (9)
where $w_{\mathrm{box}}$ and $h_{\mathrm{box}}$ represent the width and height of the smallest enclosing box that covers both the predicted and ground-truth boxes. Each exponential term decreases as the normalized offset increases due to the negative scaling factor, ensuring that $\mathrm{Dist}$ is monotonic in misalignment. Moreover, axis-wise normalization promotes scale invariance, which is crucial for handling object size variations in diverse datasets.
In summary, this angle-weighted distance penalty effectively captures both the magnitude and direction of misalignment. It provides stronger supervision in challenging cases such as small, rotated, or overlapping objects, as commonly observed in the DIOR and VisDrone datasets. Empirically, we find that incorporating this term improves detection accuracy by enhancing the localization sensitivity to angular deviation.
@@ 3.5.5 Wise Difficulty Modulation via EMA Normalization
Precise regression of small defects in UAV imagery is hindered by an imbalance in sample difficulty: abundant trivial cases dominate gradients, while naïve hard mining overfits noisy outliers and mislabels. We therefore propose Wise Difficulty Modulation (WDM)—a normalization-driven reweighting scheme that scales each sample’s gradient by an IoU-based difficulty estimate. WDM stabilizes training via an EMA-normalized difficulty distribution that adapts as accuracy improves, and uses a non-monotonic modulation to emphasize moderately hard, informative samples while down-weighting both trivial and extreme cases. This yields smoother optimization and consistently better localization, especially for small or ambiguous defects common in UAV datasets.
We quantify sample difficulty using the normalized IoU gap:
$\beta=\frac{1-\mathrm{IoU}}{\mathbb{E}[\,1-\mathrm{IoU}\,]+\varepsilon},\qquad\mathbb{E}[\,1-\mathrm{IoU}\,]\leftarrow(1-m)\,\mathbb{E}[\,1-\mathrm{IoU}\,]+m\cdot(1-\mathrm{IoU})_{\text{batch}},$ (12)
where the expectation is tracked by an exponential moving average (EMA). We use $m=10^{-2}$, which provides a stable yet responsive adaptation to phase shifts during training; in practice, a tiny constant $\varepsilon=10^{-6}$ is added to the denominator for numerical stability. Crucially, Eq. (12) operates on the raw IoU from Eq. (3) to preserve geometric interpretability. This normalization eliminates the influence of absolute IoU scale drift, allowing difficulty to be compared consistently across epochs and batch conditions.
@@
This non-monotonic weighting concentrates higher weights on samples with moderate normalized difficulty while gently down-weighting trivially easy cases and preventing domination by extreme outliers as $\beta$ increases. Empirically, such weighting increases the proportion of informative samples contributing to effective gradient updates, accelerating convergence and improving the recall of subtle, low-SNR targets without introducing instability.
Overall, Wise Difficulty Modulation via EMA Normalization offers a self-adaptive, statistically grounded mechanism to stabilize training under severe difficulty imbalance. In the context of transmission line defect datasets, it effectively improves localization precision and reduces missed detections of small and ambiguous defects by ensuring that learning focuses on the most informative samples throughout the optimization process.
@@ 3.5.6 Final Objective and Batch-wise Computation
Multiplying the geometric core (Eq. (11)) by the difficulty weight (Eq. (13)) yields the Focaler–Wise–SIoU objective. Recalling that
$L_{\mathrm{IoU}}=1-\hat{\mathrm{IoU}},\qquad L_{\mathrm{SIoU}}=L_{\mathrm{IoU}}+\tfrac{1}{2}\bigl(\mathrm{Dist}+\mathrm{Shape}\bigr),$
the final loss can be compactly written as
$\boxed{L\;=\;L_{\mathrm{SIoU}}\,\gamma(\beta),\quad\beta=\frac{1-\mathrm{IoU}}{\mathbb{E}[\,1-\mathrm{IoU}\,]+\varepsilon}}$ (14)
where $\beta$ is defined from the raw IoU as in Eq. (12), and $\gamma(\beta)$ is given in Eq. (13). The EMA-based normalization acts only through $\beta$, and no additional normalization is introduced in the final objective.
4 Results
@@ 4.2 Model Training and Evaluation Metrics
The proposed model was implemented in the PyTorch framework and built on top of the Ultralytics RT-DETR backbone. Experiments were conducted on a workstation equipped with an NVIDIA GeForce RTX 3090 (24 GB) running Ubuntu 20.04, Python 3.10.14, PyTorch 2.3.1, and CUDA 12.1. Training proceeded for 100 epochs with a batch size of 4 and an input resolution of $640\times 640$. Optimization employed AdamW with an initial learning rate of $1.0\times 10^{-4}$, a final learning-rate factor $lrf=1.0$, momentum $0.9$, and weight decay $1.0\times 10^{-4}$. We used a 2,000-iteration warm-up during which the optimizer momentum was set to $0.8$ and the bias learning rate was scaled by a factor of $0.1$.